### v1

# NEURAL EMBEDDING COMPRESSION FOR EFFICIENT MULTI-TASK EARTH OBSERVATION MODELLING

Abstract
[…]
Index Terms—  Earth Observation, Neural Compression, Foundation Models, Embeddings, Computer Vision
1 INTRODUCTION
Fig. 1: Downstream comparison: a) Raw Data Compression (RDC); b) Uniformly Quantized Embeddings (UQE); c) Neural Embedding Compression (NEC). ‘Learned’ refers to entropy coding with the distribution learned during training.
Earth observation (EO) repositories comprise some of the largest data stores globally. Due to their widespread use, these repositories experience extremely large volumes of data transfers. For example, users of the Sentinel Data Access System downloaded 78.6 PiB of data in 2022 alone [1]. The transfer of such data volumes between data producers and consumers (Fig. 1 (a)) causes substantial latency and requires significant amounts of energy and vast storage capacities.
Due to its scale, manual processing of the incoming EO data is not an option, leading to the adoption of machine learning methods, including deep learning. Recently, several EO foundation models (EO-FM) based on self-supervised learning (SSL), pre-trained on EO data sampled across the planet, have been introduced [2, 3, 4, 5, 6, 7]. The general-purpose embeddings they produce have been demonstrated as suitable for training task-specific heads for multiple downstream tasks (e.g. scene classification, semantic segmentation). The approach requires fewer annotated samples to train the task-specific model and achieves superior performance and generalization to unseen data. On the other hand, FMs consist of hundreds of millions or even billions of parameters [5, 8] and thus are computationally demanding during pre-training and at inference time.
To overcome these issues, we propose the exchange of EO-FM embeddings computed at the data provider instead of the raw data. It follows that novel data only has to be processed once by the computationally expensive encoder The resulting embeddings can be stored and reused for multiple downstream applications. The inference can be performed at the data consumer with a lightweight head.
One of the main challenges is now the efficient exchange of embeddings, which we address in this work. This is relevant as FMs do not have compression as a goal, often producing embeddings that are even larger than the original data. When transferring EO data, lossless compression is usually applied, as in Fig. 1 (a), resulting in a compressed representation of the data from which it can be reconstructed with no information loss. We refer to this method as Raw Data Compression (RDC).
One could consider applying lossless compression to the embeddings as well. For higher compression rates, lossy compression methods could be employed at the cost of some distortion. A suitable metric for this distortion is critical in elaborating a lossy compression method. To minimize the distortion, different codecs (e.g., JPEG [9] for images, MP3 [10] for audio) exist to compress data in different domains. Each of these minimizes a domain-specific metric for distortion, usually some proxy for human perception. Currently, no such codec exists in the domain of embeddings for downstream tasks.
As a simple approach, depicted in Fig. 1 (b), we could consider quantizing the embeddings to different bit levels before entropy coding. We refer to this as Uniformly Quantized Embeddings (UQE).
In this study, we propose the application of learned neural compression, which has been well studied in the domain of image and video compression [11, 12, 13, 14]. This data-driven approach allows us to learn an encoder that produces compressible multi-task embeddings from data while minimizing an appropriate distortion metric: the loss across downstream tasks. We call this approach Neural Embedding Compression (NEC) and illustrate it in Fig. 1 (c).
Our contributions in this paper are threefold: i) The discussion on EO-FM embedding sharing compared to raw data, ii) the conceptual integration of neural compression into the data pipeline, and iii) the benchmarking of the neural compression approach for various downstream applications.
2 RELATED WORK
Learned neural compression was pioneered by Balle et al. [11] with their seminal work on integrating information theoretical concepts into Deep Neural Network training for data compression. A variety of works followed proposing more powerful entropy models [12, 15] and extending its applications to video compression [14, 16, 13] and even neural network compression [17]. Recently, some works have leveraged masked auto-encoding (MAE) as a lossy compression method, dropping data at the image patch [18] or embedding level [19].
CORSA [20] trains a VQVAE [21] for the purposes of compression for satellite imagery. In their work, they briefly explore training an MLP directly on the VQVAE embeddings for a simple binary classification task, but their method does not show conclusive results. To the best of our knowledge, ours is the first work that demonstrates the effectiveness of the direct use of compressed embeddings in downstream tasks by combining FMs with neural compression. We target EO applications as input data is often sourced through queries on large repositories, whereas other application areas may rely on user input. This characteristic makes it a good fit for NEC.
3 METHOD
Fig. 2: Adapting an FM encoder for NEC.
In this section, we frame our method as the optimization of a rate-distortion objective following the framework of end-to-end learned compression established in Balle et al. [11]. Traditionally, these methods aim to transmit an input image or video using the smallest amount of bits possible (R - Rate) while minimizing the loss in quality (D - Distortion) after reconstruction by a decoder network. This trade-off is controlled by $\lambda$, resulting in an optimization problem of the form $\min\lambda R+D$.
For NEC, we are not specifically interested in reconstructing the original data. Instead, given an EO input, we aim to transmit the smallest amount of bits possible (R) so that we can perform model training or inference for various tasks, achieving the smallest loss (D) possible across them. In what follows, we detail each component of our method, summarized in Fig. 2.
Distortion Term — SSL: Our distortion term is an aggregate loss over downstream tasks which are a priori unknown. Therefore, we require a proxy for this quantity. In this work, we leverage self-supervised learning for this purpose, allowing for training on large unlabeled datasets to produce general-purpose embeddings which have shown competitive performance in many downstream tasks [22].
Specifically, we use MAE [23] with a standard vision transformer (ViT) [24] architecture. MAE masks part of the input and tasks the model with recovering it based on the unmasked input. This is achieved by minimizing the mean squared error between the masked input and its reconstruction, which we take as our distortion term. We choose MAE due to its simplicity and computational efficiency and the ViT-B architecture due to its wide use in recent EO models [5, 6, 3]. However, any loss that encourages generally useful embeddings for EO and any compatible encoder architecture can be used.
Rate Term: Balle et al. [11] establishes the general methodology for computing the rate term in the full loss. We give a short overview and direct the reader to [12, Appendix 6.1] for more details.
Given an embedding vector $y$, the minimum number of bits required to transmit $y$ is given by Shannon’s source coding theorem [25] as $-\log_{2}p(y)$, where $p$ defines the probability distribution over all embedding vectors. Our goal is to minimize the expectation of this quantity over the distribution of all embedding vectors $-\mathbb{E}\left[\log_{2}p(y)\right]$, which is precisely the entropy of this distribution. We adapt the fully factorized entropy model defined in Balle et al. [12] in order to model the probability distribution of these embeddings for MAE. Specifically, we leverage the implementation in the CompressAI [26] library.
After patching, masking, and passing through the encoder, we obtain an embedding $y$ of shape $e\times n$, where $e$ is the embedding dimension of the encoder and $n$ depends on the input shape. We quantize $y$ by rounding to the nearest integer, using uniform noise as a differentiable proxy during training [12]. To model the distribution of $y$ we assume that all of its elements are independent and additionally that all elements with the same embedding dimension (the first dimension above) are identically distributed, resulting in $p(y)=\prod_{e}\prod_{n}p_{e}(y_{e,n})$. Each $p_{e}$ is modeled using a small neural network.
SSL Compression Loss: For an input $x$ for which the model produces the (approximated) quantized embedding $y$ and reconstruction $x\prime$, the loss is given as
$Loss(x,x\prime)=\lambda\cdot\text{MAE}(x,x\prime)-\log_{2}p(y)$
with $\lambda$ as the hyperparameter controlling the trade-off.
Efficient Model Adaptation: For training, we initialize our model with the pre-trained weights from Wang et al. [3], resulting from MAE training over the MillionAid [27] dataset. We freeze all layers of this model except for the patch embedding layers of the encoder and decoder, the final encoder layer, and the first decoder layer, optimizing only $\sim$10of the total parameters.
Fine-tuning: For fine-tuning, the data consumer can train a model using the embeddings as surrogates for the raw training data. The setup is illustrated in Fig. 1 (c). We freeze the entire backbone, simulating the effect of the data consumer having access only to the embeddings.
Additionally, we experiment with the consumer having access to the first decoder layer (including the embedding layer). In this case, we unfreeze these layers and apply them as the first step of processing the embeddings after entropy decoding and before the task-specific head. The transfer of these layers’ weights is a small one-time cost that would be paid a priori, as with the probability models for the entropy codes [11].
4 EXPERIMENTS
For Neural Embedding Compression (NEC) we train on the MillionAid [27] dataset for $20$ epochs, only $1.25\%$ of the 1600 epochs used for the original pre-training. We use the same settings as Wang et al. [3] except for a reduced learning rate of $1.5\times 10^{-4}$ and 4 GPUs instead of 8, halving the total batch size to 1024. Embeddings at different compression levels are obtained by training models with different values of $\lambda$ (ranging from $10^{7}$ to $10^{11}$) but with the same architecture. We benchmark NEC against Raw Data Compression (RDC) and Uniformly Quantized Embeddings (UQE), using these embeddings to train models for two downstream EO tasks: scene classification and semantic segmentation.
In RDC, we compare against transferring the raw data, after entropy coding using the popular Zstandard [28] algorithm, at 16 and 8-bit precision, and fine-tuning the entire model, as shown in Fig. 1 (a). The 16-bit point is a useful reference despite the fact both curated datasets use 8-bit precision since many EO datasets (e.g., Sentinel 2) store data in 16-bit precision.
In UQE, we compare against transferring quantized embeddings produced by the pre-trained ViT-B network from Wang et al. [3], as shown in Fig. 1 (b). These are quantized using affine uniform quantization and entropy coded using Zstandard [28]. We show results for $32$ and $16$ bit float and $8$, $5$, $3$ and $2$ bit integer quantization.
Downstream Task Evaluation: For each task, we plot accuracy against the average size of an embedding for a sample from the validation set of the respective datasets.
1. Scene Classification: We choose the UCMerced land use dataset [29] for our scene classification task. It contains 21 classes, each with 100 images of size $256\times 256$. We use a simple architecture comprising a pooling layer followed by a single linear layer, discarding the [CLS] token. For the pooling layer, we found the ‘AttentionPoolLatent’ layer, as implemented in the ‘timm’ [30] library, to perform substantially better than other methods, such as average pooling. We train all models with a learning rate of $1.25\times 10^{-4}$ and a batch size of $32$ for 400 epochs using the AdamW [31] optimizer with weight decay of $0.05$.
2. Semantic Segmentation: The Potsdam [32] dataset is used for our semantic segmentation task. This dataset contains large images of urban areas with semantic segmentation labels for six classes. We process it as is standard [33], obtaining tiles of size $512\times 512$. Unlike Wang et al. [3], we keep the ‘clutter’ class and train with labels where the borders of semantic classes are ignored to avoid the ambiguity in labeling these. Additionally, we restrict ourselves to the RGB channels to remain compatible with the MillionAid pre-training. Most state-of-the-art models for semantic segmentation [3, 33, 8] combine embeddings from different stages of the encoder. While NEC can be extended to compress multiple levels of embeddings, we limit our investigation to the single embedding case. We use a simple decoder composed of a first convolutional layer to reduce the embedding dimension by half, followed by six residual convolutional blocks. Then, we use two PixelShuffle [34] blocks to upscale the result by a factor of four. Finally, this output is passed on one branch to a main head with two convolutional layers with $128$ channels and on another branch to an auxiliary head with one convolutional layer with $64$ channels. The loss of the auxiliary head is added to the loss of the main head with a multiplicative factor of $0.4$.
5 Results
(a) Classification Accuracy on the UC Merced dataset
(b) Overall Accuracy on the Potsdam dataset
Fig. 3: Accuracy metric vs Average Bits per Sample for Scene Classification and Semantic Segmentation
Fig. 4: Semantic segmentation on the Potsdam dataset. The mean size of the embedding produced by the method over the validation set is shown in KB.
1. Scene Classification: Fig. 3(a) plots the results for the UCM Dataset. For the high-performance case, we observe that UQE performs well until $5$ bit quantization and is only slightly outperformed by NEC. For higher compression levels, this does not hold. NEC is able to achieve $94\%$ accuracy at an average embedding size of $0.47$KB, a reduction of only $5\%$ in accuracy for a size reduction of $\sim$700700700700x, greatly outperforming 3 and 2-bit UQE. The benefit of transmitting the first layer of the decoder is also apparent, particularly as the compression level increases.
2. Semantic Segmentation: Results are shown in Fig. 3(b). NEC scales better with compression, achieving more than a $100$x reduction in size with only $5\%$ reduced overall accuracy, compared to a $12\%$ drop in accuracy for comparable compression using UQE. Fig. 4 shows a visual comparison of the inference outputs. For a smaller bit budget on this difficult image, with many separate occurrences of classes to identify, NEC produces more accurate segmentation masks.
In both tasks, we demonstrate the practical viability of our method, with only a small performance gap between RDC and NEC , despite not fine-tuning the encoder.
6 CONCLUSION
In this work, we introduce NEC, a framework for downstream training and inference for EO tasks based on directly transmitting compressed embeddings. We show how any FM can be adapted to produce compressed embeddings of different fidelities using a single architecture and with minimal compute. Our evaluation concludes with a data transfer reduction of one to three orders of magnitude with only limited degradation in performance.
We argue the utility of our work from a sustainability point of view, enabling training and inference with limited bandwidth or computational resources and reducing data transfer, storage requirements, and their associated energy costs. Furthermore, we see potential in the long-term storage of compressed embeddings produced from large repositories of EO data (e.g., Landsat, Sentinel) for future analysis.